OpenStudy (lukecrayonz):

Find all solutions in the interval [0, 2π). sec^2 x - 2 = tan^2 x HELP! :o

OpenStudy (lukecrayonz):

I'm pretty sure the answer is no solution..

OpenStudy (raden):

there is no solution... maybe, the typo there

OpenStudy (anonymous):

\[\sec^2 x - 2 = \tan^2 x \]\[\sec^2 x - 1 = \tan^2 x +1\]\[\sec^2x=\tan^2x+1\]\[\sec^2x-1=\sec^2x\]\[-1 \not= 0\]No Solution

OpenStudy (lukecrayonz):

okay thank you so much! can you help me with one more question? @ChmE

OpenStudy (anonymous):

sure

OpenStudy (lukecrayonz):

A high-altitude spherical weather balloon expands as it rises due to the drop in atmospheric pressure. Suppose that the radius r increases at the rate of 0.16 inches per second and that r = 38 inches at time t = 0. Determine the equation that models the volume V of the balloon at time t and find the volume when t = 280 seconds.

OpenStudy (anonymous):

btw here is a good reference site for these types of problems http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

idk sry. I can't do word problems

OpenStudy (lukecrayonz):

I have another then! @ChmE

OpenStudy (lukecrayonz):

Find the inverse of 3√(x/8) -4

OpenStudy (anonymous):

\[3\sqrt{{x \over 8} -4}\]is this the question

OpenStudy (lukecrayonz):

yes! the three is the little number

OpenStudy (lukecrayonz):

@sirm3d help with the inverse question please!

OpenStudy (sirm3d):

starting with x, divide by 8, subtract by 4, take the cube root Inverse: starting with x, cube it, add 4, then multiply by 8

OpenStudy (sirm3d):

it's like socks before shoes. the inverse: shoes before socks

OpenStudy (sirm3d):

the inverse of the function is \[\large (x^3+4)(8)\]

OpenStudy (lukecrayonz):

thank you! you're a lifesaver

OpenStudy (sirm3d):

^^